Identify key elements of the Ball and Brown (1968) methodology and discuss the contribution of the study to subsequent research into the association between earnings and equity returns. Discuss the contribution of the Ball and Brown (1968) methodology to our understanding of the relationship between earnings and equity returns.
Answer #1

Relationship between earnings and stock returns

empirical evidence from the Greek Capital Market

earning and return specification

price model Pt= a+ bXt+et

return model :Pt/Pt-1=a+bXt/Pt-1+et

differenced model:^Pt/Pt-1=a+bXt/Pt-1+et/Xt +b+et

Pt is ex- dividend price

t financial year

X is earning per share excluding taxes and discontinued operations for period t , a and b are intercept and the earning coefficient respectively, e is the error term.